OpenStudy (anonymous):

(3 x 10²) m + (5 x 10²) m

OpenStudy (turingtest):

300m+500m=80000m=8x10^4m

OpenStudy (anonymous):

the answer is 8 times 10² but i want to know why it isn't to the fourth

OpenStudy (turingtest):

oh man i messed up 3x10^2=300 5x10^2=500+ =800=8x10^2m

OpenStudy (anonymous):

Ohh so you would not add the two ^2 to make it 4?

OpenStudy (lalaly):

\[\huge{10^2 \times 10^2 = 10^{2+2}}\]\[\huge{10^2+10^2=2 \times 10^2}\]

OpenStudy (lalaly):

so in this case =8 x 10^2

OpenStudy (anonymous):

\[8 \times 10^{-5} + 8 \times 10^{-5}\]

OpenStudy (lalaly):

same thing add 8+8=16 so 16 x 10^-5

OpenStudy (anonymous):

the answer is 1.2 times 10^-4

OpenStudy (anonymous):

i don't know how to get it though

OpenStudy (lalaly):

1.6 x 10^-4

OpenStudy (anonymous):

no.. 1.2 times 10^-4

OpenStudy (lalaly):

its not 1.2 its 1.6

OpenStudy (anonymous):

how do you get 1.6 though

OpenStudy (lalaly):

8+8=16 so 16 x 10^-5 = 1.6x 10^-4

OpenStudy (anonymous):

it's 8 +4

OpenStudy (lalaly):

\[8 \times 10^{-5} + 4 \times 10^{-5} ?\]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Sorrry i messed up up there^^

OpenStudy (lalaly):

ok since both 10's hav the same exponent u add 8+4=12 so the answer is 12x10^-5 = 1.2x10^-4

OpenStudy (lalaly):

A=10^2*m so it will be: 3A+5A (3+5)A same thing wit the other one

OpenStudy (anonymous):

thank you so much.